Solve the inequality √(x2-2x-8) ≤ -x+2


1
Expert's answer
2021-11-01T10:59:50-0400

"\sqrt {x^2-2x-8}\leq -x+2" "...(1)"

We know that, if "\sqrt {f(x)}\leq g(x)" then "g(x) \geq0"

So, "-x+2\geq0\Rightarrow x\leq2 \ ...(2)"

As "-x+2\geq0"

So, squaring both sides in equation (1), we get

"x^2-2x-8\leq(-x+2)^2
\\\Rightarrow x^2-2x-8\leq x^2+4-4x
\\\Rightarrow 2x\leq12
 \\\Rightarrow x\leq6 \ ...(3)"

Also "x^2-2x-8\geq 0" as value inside square root can not be negative.

"\\\Rightarrow (x+2)(x-4)\geq0"



From the above, we can say that

"x\leq-2" or "x\geq4 \ ...(4)"

From equation (2), (3) and (4)



From the graph, we can say that the common solution is

"x\leq-2"